What is the quantity of heat required to change 0.02 kg of ice at 0 oC to water at 60 oC ?

[ specific heat capacity of water = 4.2 x 10^3 J kg^-1 o C^-1 ]

[ specific latent heat of fusion of ice = 3.34 x 10^5 J kg^-1 ]


1
Expert's answer
2021-07-15T10:51:27-0400

Let's first find the quantity of heat required to convert "0.02\ kg" of ice at "0^{\circ}C" to "0.02\ kg" of water at "0^{\circ}C":


"Q_1=mL_f=0.02\ kg\cdot3.34\cdot10^5\ \dfrac{J}{kg}=6680\ J."

Then, we can find the quantity of heat required to convert "0.02\ kg" of water at "0^{\circ}C" to "0.02\ kg" of water at "60^{\circ}C":


"Q_2=mc\Delta T=0.02\ kg\cdot4200\ \dfrac{J}{kg\cdot \!^{\circ}C}\cdot60^{\circ}C=5040\ J."

Finally, we can find the total amount of heat required to convert 0.02 kg of ice at "0^{\circ}C" to water at "60^{\circ}C":


"Q=Q_1+Q_2=6680\ J+5040\ J=11720\ J."